OpenStudy (anonymous):

write 1 and 2/3 as an improper fraction?

OpenStudy (one098):

Multiply 3 times 1 then add 2.

OpenStudy (anonymous):

solve it then put it in a improper fraction?

OpenStudy (texaschic101):

1 2/3 to turn a mixed number into an improper fraction... take the whole number (1) and multiply it by the denominator (3) giving us 1 x 3 = 3...then add that to the numerator (2), giving us 3 + 2 = 5...now put that over the original denominator (3)...giving us 5/3. so basically, its : (1 x 3 + 2) / 3 = 5/3

OpenStudy (anonymous):

ohh lol read that wrong